OpenStudy (anonymous):

SOMEONE PLEASE HELP ME! Which equation is equivalent to y-6 = -12(x+4)? A. y = -6x - 48 B. y = 6x - 48 C. y = -12x - 42 D. y = -12x - 54 Which point is located on the line represented by the equation y + 4 = -5(x-3) A. (-4, -5) B. (-5, -4) C. (3, -4) D. (-3, 4)

OpenStudy (anonymous):

Anyone, please? I'll do anything!

OpenStudy (anonymous):

D

OpenStudy (anonymous):

D is the first one

OpenStudy (anonymous):

What is the second one?

OpenStudy (anonymous):

I dont remeber how to do that type

OpenStudy (danjs):

the first you just have to rearrange things to solve for y = ... y - 6 = -12x - 48 y = -12x -42

OpenStudy (danjs):

a point (x,y) is on the given line, if it makes the equation true recall however, that is the 'point-slope ' form of a line y - y1 = m*(x - x1)

OpenStudy (danjs):

the point with terms, ( x1, y1) may be in the answers

OpenStudy (danjs):

y + 4 = -5(x-3) y -(-4) = -5*(x - 3) y1 = -4 x1 = 3 (3,-4) is one of the choices,